OpenStudy (anonymous):

y + 2x = -1 3y - x = 4 A. (-1,1) B. (1,1) C. no solution

OpenStudy (anonymous):

y + 2x = -1 3y - x = 4 Rearrange 2x+y = -1 -x+3y = 4

OpenStudy (anonymous):

so C? sorry idk

OpenStudy (anonymous):

@ChiefArnav

OpenStudy (anonymous):

2x+y = -1 multiply yhis by 2 *this

OpenStudy (anonymous):

multiply -1 by 2?

OpenStudy (anonymous):

or2x by 2?

OpenStudy (anonymous):

2x+y = -1 each term by 2 i.e. 2x*2 + 2*y = -1*2

OpenStudy (anonymous):

4x+?=2 or -2 sorry i suck at math :/

OpenStudy (anonymous):

4x+2y = -2

OpenStudy (anonymous):

oh ok almost had it lol

OpenStudy (anonymous):

2x+y = -1 -x+3y = 4 Take 2x+y=-1 y = -1-2x Use this value of y in -x+3y = 4 i.e. -x+3(-1-2x) = 4

OpenStudy (anonymous):

-x - 3 - 6x = 4 -x-6x = 4+3 -7x = 7 x = 7/-7 x = -1

OpenStudy (anonymous):

Now, y = -1-2x use x = -1 here y = -1-2(-1) y = -1+2 y = 1 Therefore,A

OpenStudy (anonymous):

thnx :)